# New black hole mergers in the LIGO–Virgo O3 data from a gravitational wave search including higher-order harmonics

### v1

@@ V.2 What is the gain/loss in sensitivity due to including HM in the search?
We also find that in our new search, the significance of a few previously reported marginal events is increased significantly (GW190711_030756, GW200216_220804). Especially for the case of GW190711_030756, we suspect that the increase in sensitivity is due to addition of HM, but to test this we plan to do an injection study in the future.
The downside of using a prior distribution on is that it is dependent on our choice of astrophysical prior (which is currently not precisely known). Especially, the mass ratio distribution is not very well constrained and affects our prior for . Using a different distribution for than the true astrophysical distribution can indeed lead to some loss of sensitivity. We chose a broad prior (flat in ) in our search, with the motivation that such a prior can be easily reweighted later in a hierarchical population analysis.
Even though including HM can cause some loss in sensitivity for near-threshold equal mass systems, note that detecting even a few events with asymmetric-mass ratios can inform our population estimates much more than detecting a slightly larger number of equal-mass systems. It is worth drawing an analogy here to the case of spin distribution of templates used in -only searches. Neglecting templates with negative indeed improves the search sensitivity to near-threshold systems, but finding even a few systems can greatly help us distinguish between progenitor channels of BHs. We thus argue the goal of a search should not solely be to maximize the total number of detections, but to holistically cover most of the interesting regions of the binary parameter space. A similar argument also holds for a precessing search, where detecting even a few events with significantly misaligned spins can shed more light on the formation channels of binaries as compared to detecting a larger number of near-threshold aligned/zero spin mergers.
@@ VI Conclusions
Nearly all gravitational searches performed till now only include the quadrupole mode of the full GW waveform predicted by general relativity. In our companion papers Wadekar et al. (2023a, b), we develop new efficient methods to introduce higher-order modes (HM) in our search pipeline (section II). This helped us perform the first search over the full binary black hole parameter space (i.e., over all binary inclinations) with including HM in our templates. Compared to our previous pipeline, we introduced additional improvements to remove/downweight noise transients (especially at low frequencies), which further improved our sensitivity to high mass and high redshift systems (section II.3).
@@
We searched over public data from the third LIGO–Virgo observing run and found 14 new candidate events with (see Table 1 and section III). Some of the new candidate events have several interesting astrophysical properties (section IV) such as: () masses in the IMBH range and pair instability mass gap, () high redshifts, () preference for positive effective spins, () asymmetric mass ratios.
@@
By simply summing over the complements of values of our new candidate events from Table 1, we expect that roughly 4 of the 14 new events are noise transients instead of being astrophysical signals. However, we note that estimates of and source parameters depend on the choice of prior, and we discuss this dependence in section V555We provide our posterior samples at https://github.com/JayWadekar/GW_higher_harmonics_search, which can be used to reweight our results for a different astrophysical prior..
@@
We also compared our pipeline’s results for the events published in previous IAS catalogs (Tables 2 & 6), and in the GWTC-3 and 4-OGC catalogs (Table 3). We broadly recovered the high-significance events, except some events which were either vetoed by our pipeline or fell below our trigger collection SNR threshold, which we discuss in detail in section III.3 (we also mention some possible future improvements in this regard).
@@
Our methodology for efficiently including realistic effects like HM in GW templates will continue to be useful for high-precision searches in upcoming and future gravitational wave experiments Fairhurst et al. (2023); Maggiore et al. (2020); Evans et al. (2023); Kawamura et al. (2020); Amaro-Seoane et al. (2017); Akutsu et al. (2021). One particular example in which HM will be especially useful is when searching for extreme-mass-ratio inspirals in LISA Zhang et al. (2023); Baibhav et al. (2020).
VI.1 Future work
We plan to inject simulated gravitational-wave signals in the data and compare the fraction of injections recovered by our pipeline. The injections will help us determine the gain/loss in volume-time sensitivity in different regions of the parameter space due to including HM in our search. The injections will also help us gauge the level of improvement due to the additional glitch treatment methods that we described in section II.3. We plan to use the injection parameters from Abbott et al. (2021), which will help us perform an apples-to-apples comparison of our sensitivity with other pipelines. Having injections will also help quantify the selection function of our current pipeline and help in population inference studies. The new events reported here have moderate IFARs (the best event has yr), thus a methodology to combine confident and marginal events in population inference will be required Roulet et al. (2020). We expect our sensitivity to be higher than other searches for asymmetric mass ratio events due to addition of HM. Our sensitivity could also be higher for high-mass and high-redshift regions due to using improved glitch treatment methods and adding HM. Note that, in certain regions of the parameter space, having higher sensitivity can improve population estimates even if we have null/marginal detections. We note some interesting population inference questions in section IV.
Currently, nearly all searches also ignore the effects of precession or eccentricity in their waveforms. Some of the challenges for including these effects are similar to the ones for HM: significant increase in the template bank size and loss of sensitivity due to collecting additional background. We hope to develop methods similar to the ones implemented in our search to help tackle those searches in the future.
@@ Acknowledgements.
We thank Horng Sheng Chia, Ajith Parameswaran, Will Farr, Vishal Baibhav, Katerina Chatziioannou, Vicky Kalogera, Sylvia Biscoveanu, Amanda Farah, Daniel Holz, Salvatore Vitale, Tom Edwards, Zoehyr Doctor, Suvodip Mukherjee and Mukesh Singh for helpful discussions. DW gratefully acknowledges support from the Friends of the Institute for Advanced Study Membership and the Keck foundation. TV acknowledges support from NSF grants 2012086 and 2309360, the Alfred P. Sloan Foundation through grant number FG-2023-20470, the BSF through award number 2022136, and the Hellman Family Faculty Fellowship. MZ is supported by NSF 2209991 and NSF-BSF 2207583. BZ is supported by the Israel Science Foundation, NSF-BSF and by a research grant from the Willner Family Leadership Institute for the Weizmann Institute of Science. This research was also supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. We also thank ICTS-TIFR for their hospitality during the completion of a part of this work. This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.Appendix A Additional plots and tables

### v2

@@ V.3 Population study
We also find that in our new search, the significance of a few previously reported marginal events is increased significantly. Especially for the case of GW190711_030756, the increase in sensitivity is due to addition of HM. This will also be important in population level studies. Currently the LVK collaboration uses events with IFAR 1 year in their population studies. Due to the improved significance of our pipeline, more events can now be used in the IFAR 1 analysis
@@ VI Conclusions
Nearly all the previous gravitational-wave searches only include the quadrupole mode of the full GW waveform predicted by general relativity. In our companion papers Wadekar et al. (2023, 2024), we develop new efficient methods to introduce higher-order modes (HM) in our search pipeline (section II). This helped us perform the first search over the full binary black hole parameter space with including HM in our templates. Compared to our previous pipeline, we introduced additional improvements to remove/downweight noise transients (especially at low frequencies), which further improved our sensitivity to high mass and high redshift systems (section II.3).
@@
We searched over public data from the third LIGO–Virgo observing run and found 11 new candidate events with (see Table 1 and section III). Some of the new candidate events have several interesting astrophysical properties (section IV) such as: () masses in the IMBH range and pair instability mass gap, () high redshifts, () non-zero effective spins, () asymmetric mass ratios.
@@
By simply summing over the complements of values of our new candidate events from Table 1, we expect that roughly 3 of the 11 new events are noise transients instead of being astrophysical signals. However, we note that estimates of and source parameters depend on the choice of prior, and we discuss this dependence in section V666See https://github.com/JayWadekar/GW_higher_harmonics_search for obtaining our posterior samples and the selection function of our pipeline..
@@
We also compared our pipeline’s results for the events published in previous IAS catalogs (Tables 2 & 6), and in the GWTC-3 and 4-OGC catalogs (Table LABEL:tab:lvc_events). We broadly recovered the high-significance events, except some events which were either vetoed by our pipeline, which we discuss in detail in section III.3.
@@
Our methodology for efficiently including realistic effects like HM in GW templates will continue to be useful for high-precision searches in upcoming and future gravitational wave experiments Ng et al. (2023); Fairhurst et al. (2023); Maggiore et al. (2020); Evans et al. (2023); Kawamura et al. (2020); Amaro-Seoane et al. (2017); Akutsu et al. (2021). One particular example in which HM will be especially useful is when searching for intermediate to extreme mass-ratio inspirals Arun et al. (2007a, b); Porter and Cornish (2008); Berti et al. (2006); Baibhav et al. (2020); Pitte et al. (2023); Amaro-Seoane et al. (2023); Gair et al. (2008); we report results from such a search in a separate follow-up paper Ho-Yeuk Cheung et al. (2025).
Note added– While this paper was on arXiv, a confident high-mass detection (GW231123) was reported by the LVK collaboration Abac et al. (2025) in the O4a data. Incorporating this detection into the search astrophysical prior could make our high-mass detections () more significant.
@@ Acknowledgements.
We thank Horng Sheng Chia, Ajith Parameswaran, Will Farr, Vishal Baibhav, Katerina Chatziioannou, Vicky Kalogera, Sylvia Biscoveanu, Amanda Farah, Daniel Holz, Salvatore Vitale, Tom Edwards, Zoheyr Doctor, Suvodip Mukherjee and Mukesh Singh for helpful discussions. DW gratefully acknowledges support from the Friends of the Institute for Advanced Study Membership and the Keck foundation. JR acknowledges support from the Sherman Fairchild Foundation. TV acknowledges support from NSF grants 2012086 and 2309360, the Alfred P. Sloan Foundation through grant number FG-2023-20470, the BSF through award number 2022136, and the Hellman Family Faculty Fellowship. MZ is supported by NSF 2209991 and NSF-BSF 2207583. BZ is supported by the Israel Science Foundation, NSF-BSF and by a research grant from the Willner Family Leadership Institute for the Weizmann Institute of Science. This research was also supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. We also thank ICTS-TIFR for their hospitality during the completion of a part of this work. This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.Appendix A Changes since previous arXiv version
The previous version of our paper on the arXiv ([v1] 11 Dec 2023) included the list of events from an older version of our pipeline. Since that point, we have made a number of improvements to the pipeline. Furthermore, unlike the older version, the newer version of our IAS-HM pipeline has gone through an extensive injection recovery analysis to test its senstitivity (see section V.1). We list the details of other improvements below.
We implemented a few extra noise reduction algorithms like TIER (see section II.3.1). We see an overall improvement in sensitivity due to it Wadekar et al. (a). We also however see both positive and negative fluctuations in IFAR and values of the near-threshold new candidate events. This is expected, as some of the near-threshold candidate events were probabilistically likely to be glitches and the TIER algorithm is able to downweight those.
We fixed an issue with relative binning for template optimization in our bank BBH-1. This had earlier led to some of the LVK mergers being spuriously vetoed for this bank. We have now fixed the issue.
In the previous version of our catalog, some of the fairly high SNR events in GWTC-3 with imprints of precession were vetoed by our pipeline. We added a ML-based veto overriding procedure (see section II.3.3) to override the veto for these events and we also find an improvement in the pipeline sensitivity due to this procedure.
In section II.4, we discussed using an astrophysical prior for the templates. We improved the astrophysical prior applied to the templates using normalizing flows Wadekar et al. (b), superseding the earlier kernel density estimation based method given in Appendix C of O22. We also improved our coherent marginalized statistic using normalizing flows (further details will be provided in a separate paper Wadekar et al. (b)).
Appendix B Additional plots and tables